OpenStudy (anonymous):

equations and inequalities 5x-4=2x+6

OpenStudy (anonymous):

5x-4=2x+6 subtract 2x to both sides 3x -4 = 6 add 4 to both sides 3x= 10 divide both by 3 x = 10/3